OpenStudy (anonymous):

A cylindrical container has a volume of 36 cm³. Another container made out of the same material is similar to the first container with a height that is three-fourths the height of the first container. To the nearest tenth of a cm³, what is the volume of the second container?

OpenStudy (cggurumanjunath):

V cylinder = pi * r^2*h

OpenStudy (cggurumanjunath):

v =36 cm^3

OpenStudy (cggurumanjunath):

v2=pi*r^2 *(3/4h)

OpenStudy (cggurumanjunath):

v1=pi*r^2*h =36 v2= pi *r ^2 *(3h/4) =36*3/4=9*3=27 cm^3